The average cholesterol content of a 25 canned goods is 215 miligrams,and the standard deviation is 15 miligrams assume that the variable is normally distributed.if sample of 25 canned goods is selected what is the probability that the mean of the sample will be greater than 220 miligrams


1
Expert's answer
2021-05-21T07:27:38-0400

"n=25,\mu=215,\sigma=15"


Required Probability-


"P(X>220)=P(Z>\dfrac{220-215}{\frac{15}{\sqrt{25}}}" )


"=P(Z>1.66)=0.04779"